OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (karon86):

Use the formula y-y1=m(x-x1)

OpenStudy (karon86):

m=slope (x1,y1)

OpenStudy (karon86):

(x1,y1) are the points that passes through the slope

OpenStudy (anx):

y+4 = -3(x-3) distrubute the -3 and add 4 to both sides y+4 = -3x -9 +4 +4 y= -3x-5

OpenStudy (karon86):

that would be the linear equation but the equation is y+4=-3(x-3)

OpenStudy (karon86):

the equation in point-slope form is y+4=-3(x-3). You do not have to go any further unless the question asks for the Slope intercept form so basically you have the right answer

OpenStudy (anonymous):

Thank you Everyone for helping me